Use the given zero to find the remaining zeros of the function.

f(x)=x^3-2x^2+36x-72, zero: 6i
1
Expert's answer
2019-12-06T09:18:17-0500

f (x) = x3 - 2x2 + 36x - 72 = 0, x1 = 6i

Since there is no i in the equation, suppose that the root is -6i

f (x) = x3 - 2x2 + 36x - 72 = 216i + 72 - 216i -72 = 0

The assumption turned out to be true: x2 = -6i

(x - 6i)*(x + 6i) = x2 - 6ix + 6ix + 36 = x2 + 36

Divide the function f(x) by the product x2+36 of the roots

x3 -2x2 + 36x - 72 | x2+36

x3 + 36x | x - 2

________

-2x2 -72

-2x2 -72

__

0

x3 - 2x2 + 36x - 72 = (x - 6i)*(x + 6i)*(x - 2) = 0

x-2= 0

x3 = 2

Answer:

x1 = 6i; x2 = -6i; x3 = 2.